OpenStudy (anonymous):

When a number is divided by 13, the remainder is 11. When the same number is divided by 17, the remainder is 9.. What is the number?? Options: a)339 b) 349 c) 369 d) Data Inadequate I only need the solution to this problem.. Thanks if anyone can help me with this..

OpenStudy (shubhamsrg):

if no. =x then x = 13n + 11 and x = 17m + 9 3 variables ->2 eqn.. i'd say d)

OpenStudy (shubhamsrg):

17m - 13n = 2 is there any way for integral solutions ? or just hit and trial ? :| maybe the ans is not d)

OpenStudy (anonymous):

There is an answer of this question... Answer is not d)..

OpenStudy (anonymous):

the answer is b because i tried it!

OpenStudy (anonymous):

349

OpenStudy (anonymous):

I need the solution @eliassaab ..

OpenStudy (anonymous):

Can you please explain??

OpenStudy (anonymous):

Do not try from the options.. Suppose you don't know the answers, then tell me how to get to the answer by solving it logically..

OpenStudy (anonymous):

349=26(13)+11 349 =20(17)+9

OpenStudy (anonymous):

What if I did not the give the choices??? Suppose there is no choice, then is their a method to get the answer by proper solution??

OpenStudy (anonymous):

141, 349, 557, 765, 973, etc. i just played around and try things :s

OpenStudy (anonymous):

answer is 221m-93 m is 1,2,3,...

OpenStudy (anonymous):

let x=13k+11=17y+9

OpenStudy (anonymous):

From here: k = (17y - 2)13..

OpenStudy (anonymous):

very well

OpenStudy (anonymous):

Find the value of y for which there will be no remainder...

OpenStudy (anonymous):

then k=y+(4y-2)/13

OpenStudy (anonymous):

How you did this?

OpenStudy (anonymous):

k=(13y+4y-2)/13

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

How 4y come??

OpenStudy (anonymous):

17y=13y+4y

OpenStudy (anonymous):

Here is a list of possible answers less than a million {128, 349, 570, 791, 1012, 1233, 1454, 1675, 1896, 2117, 2338, 2559, \ 2780, 3001, 3222, 3443, 3664, 3885, 4106, 4327, 4548, 4769, 4990, \ 5211, 5432, 5653, 5874, 6095, 6316, 6537, 6758, 6979, 7200, 7421, \ 7642, 7863, 8084, 8305, 8526, 8747, 8968, 9189, 9410, 9631, 9852, \ 10073, 10294, 10515, 10736, 10957, 11178, 11399, 11620, 11841, 12062, \ 12283, 12504, 12725, 12946, 13167, 13388, 13609, 13830, 14051, 14272, \ 14493, 14714, 14935, 15156, 15377, 15598, 15819, 16040, 16261, 16482, \ 16703, 16924, 17145, 17366, 17587, 17808, 18029, 18250, 18471, 18692, \ 18913, 19134, 19355, 19576, 19797, 20018, 20239, 20460, 20681, 20902, \ 21123, 21344, 21565, 21786, 22007, 22228, 22449, 22670, 22891, 23112, \ 23333, 23554, 23775, 23996, 24217, 24438, 24659, 24880, 25101, 25322, \ 25543, 25764, 25985, 26206, 26427, 26648, 26869, 27090, 27311, 27532, \ 27753, 27974, 28195, 28416, 28637, 28858, 29079, 29300, 29521, 29742, \ 29963, 30184, 30405, 30626, 30847, 31068, 31289, 31510, 31731, 31952, \ 32173, 32394, 32615, 32836, 33057, 33278, 33499, 33720, 33941, 34162, \ 34383, 34604, 34825, 35046, 35267, 35488, 35709, 35930, 36151, 36372, \ 36593, 36814, 37035, 37256, 37477, 37698, 37919, 38140, 38361, 38582, \ 38803, 39024, 39245, 39466, 39687, 39908, 40129, 40350, 40571, 40792, \ 41013, 41234, 41455, 41676, 41897, 42118, 42339, 42560, 42781, 43002, \ 43223, 43444, 43665, 43886, 44107, 44328, 44549, 44770, 44991, 45212, \ 45433, 45654, 45875, 46096, 46317, 46538, 46759, 46980, 47201, 47422, \ 47643, 47864, 48085, 48306, 48527, 48748, 48969, 49190, 49411, 49632, \ 49853, 50074, 50295, 50516, 50737, 50958, 51179, 51400, 51621, 51842, \ 52063, 52284, 52505, 52726, 52947, 53168, 53389, 53610, 53831, 54052, \ 54273, 54494, 54715, 54936, 55157, 55378, 55599, 55820, 56041, 56262, \ 56483, 56704, 56925, 57146, 57367, 57588, 57809, 58030, 58251, 58472, \ 58693, 58914, 59135, 59356, 59577, 59798, 60019, 60240, 60461, 60682, \ 60903, 61124, 61345, 61566, 61787, 62008, 62229, 62450, 62671, 62892, \ 63113, 63334, 63555, 63776, 63997, 64218, 64439, 64660, 64881, 65102, \ 65323, 65544, 65765, 65986, 66207, 66428, 66649, 66870, 67091, 67312, \ 67533, 67754, 67975, 68196, 68417, 68638, 68859, 69080, 69301, 69522, \ 69743, 69964, 70185, 70406, 70627, 70848, 71069, 71290, 71511, 71732, \ 71953, 72174, 72395, 72616, 72837, 73058, 73279, 73500, 73721, 73942, \ 74163, 74384, 74605, 74826, 75047, 75268, 75489, 75710, 75931, 76152, \ 76373, 76594, 76815, 77036, 77257, 77478, 77699, 77920, 78141, 78362, \ 78583, 78804, 79025, 79246, 79467, 79688, 79909, 80130, 80351, 80572, \ 80793, 81014, 81235, 81456, 81677, 81898, 82119, 82340, 82561, 82782, \ 83003, 83224, 83445, 83666, 83887, 84108, 84329, 84550, 84771, 84992, \ 85213, 85434, 85655, 85876, 86097, 86318, 86539, 86760, 86981, 87202, \ 87423, 87644, 87865, 88086, 88307, 88528, 88749, 88970, 89191, 89412, \ 89633, 89854, 90075, 90296, 90517, 90738, 90959, 91180, 91401, 91622, \ 91843, 92064, 92285, 92506, 92727, 92948, 93169, 93390, 93611, 93832, \ 94053, 94274, 94495, 94716, 94937, 95158, 95379, 95600, 95821, 96042, \ 96263, 96484, 96705, 96926, 97147, 97368, 97589, 97810, 98031, 98252, \ 98473, 98694, 98915, 99136, 99357, 99578, 99799}

OpenStudy (anonymous):

@waterineyes is that right?

OpenStudy (anonymous):

What people are trying to explain to you that 2 equations with three unknown might have an infinite number of solutions.

OpenStudy (shubhamsrg):

maybe i just got it..using @eliassaab work.. first no. following the series is 128 LCM(17,13) =221 thus every no is in the form of 221p + 128 p is any integer.. so you might arrive to ans thus..

OpenStudy (anonymous):

Ya I got, further.??.

OpenStudy (shubhamsrg):

since all options are 300 something,,so you can easily judge that p=1// following which you have the ans : 349

OpenStudy (anonymous):

sorry i lost my connection

OpenStudy (anonymous):

I am getting somewhat.. I have tried one more method...

OpenStudy (anonymous):

well then \[k=y+\frac{4y-2}{13} \\ k \ is \ an \ integer \ then \ \frac{4y-2}{13} \ \ must \ be \ an \ integer \\ then \ 4y-2=13t\]

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

x = 13p + 11 x = 17q + 9 Just equate: 13p + 11 = 17q + 9 p = (17q - 2)/13 Now at value of q we get p a whole number?? Let us try for q = 7 p = (119 - 2)/13 = 9 So, q = 7 and p = 9 So, the number is : 13p + 11 = 128 As this is not in the choice, so let us try once more.. Put q = 20, p = (340 - 2)/13 = 26 So, q = 20 and p = 26 So other number will be: 13p + 11 = 338 + 11 = 349...

OpenStudy (anonymous):

Thanks guys.... Thanks @mukushla thanks @eliassaab and thanks @shubhamsrg ..

terenzreignz (terenzreignz):

So it all boils down to finding the first number that satisfies the conditions (be it by trial and error) and just keeping on adding the least common multiple to that until the answer is one which is in the choices, right?

OpenStudy (anonymous):

\[4y-2=13t \ \ \ \ then \ \ 4y=13t+2 \ \ \ \\ y=3t+\frac{t+2}{4} \ \ y \ is \ an \ integer \ \ \ so \ \ \ \ \ t+2=4m \\ that \ \ results\ \ in \\ t=4m-2 \\ y=13m-6 \\ and \ \ finally \\ x=17y+9=221m-93 \ \] put m=2 then x=349

OpenStudy (anonymous):

@waterineyes your welcome

OpenStudy (anonymous):

Go thru the Options.... if not any of three means opntion d is correct

OpenStudy (anonymous):

Thanks one again @mukushla

OpenStudy (anonymous):

Actually this I am seeing today..